# Beyond Freezing: Sparse Tuning Enhances Plasticity in Continual Learning with Pre-Trained Models

### v1

@@ Appendix E Experiments on Domain-Incremental Learning
To further validate the general applicability of our method, we also conduct experiments on domain-incremental learning using the DomainNet dataset Peng et al. [2019], which is a large-scale benchmark dataset containing images from 345 categories across six diverse visual domains. Table A2 compares the proposed SimpleCIL+MIST method with several baseline approaches, including L2P, Adam, and SimpleCIL. As shown, SimpleCIL+MIST achieves the highest final accuracy (53.5%), indicating its effectiveness in DIL scenarios as well.
@@
Table A2: Comparison of different methods for domain-incremental learning evaluated on the DomainNet dataset. The proposed SimpleCIL+MIST achieves the highest final accuracy.

### v2

@@ Appendix G Experiments on Domain-Incremental Learning
To further validate the general applicability of our method, we also conduct experiments on domain-incremental learning using the DomainNet dataset (domainnet), which is a large-scale benchmark dataset containing images from 345 categories across six diverse visual domains. Table 7 compares the proposed SimpleCIL+MIST method with several baseline approaches, including L2P, Adam, and SimpleCIL. As shown, SimpleCIL+MIST achieves the highest final accuracy (53.5%), indicating its effectiveness in DIL scenarios as well.
@@
Table 7: Comparison of different methods for domain-incremental learning evaluated on the DomainNet dataset. The proposed SimpleCIL+MIST achieves the highest final accuracy.